# Is Translation All You Need? A Study on Solving Multilingual Tasks with Large Language Models

### v2

@@ Abstract
Large language models (LLMs) have demonstrated multilingual capabilities; yet, they are mostly English-centric due to the imbalanced training corpora. Existing works leverage this phenomenon to improve their multilingual performances through translation, primarily on natural language processing (NLP) tasks. This work extends the evaluation from NLP tasks to real user queries and from English-centric LLMs to non-English-centric LLMs. While translation into English can help improve the performance of multilingual NLP tasks for English-centric LLMs, it may not be optimal for all scenarios. For culture-related tasks that need deep language understanding, prompting in the native language tends to be more promising as it better captures the nuances of culture and language. Our experiments reveal varied behaviors among different LLMs and tasks in the multilingual context. Therefore, we advocate for more comprehensive multilingual evaluation and more efforts toward developing multilingual LLMs beyond English-centric ones.
@@ A Study on Solving Multilingual Tasks with Large Language Models
Chaoqun Liuthanks: Chaoqun Liu is under the Joint PhD Program between DAMO Academy and Nanyang Technological University.   12  Wenxuan Zhangthanks: Wenxuan Zhang is the corresponding author.  23  Yiran Zhao24  Anh Tuan Luu1  Lidong Bing23 1Nanyang Technological University, Singapore; 2DAMO Academy, Alibaba Group, Singapore 3Hupan Lab, 310023, Hangzhou, China; 4National University of Singapore {chaoqun.liu,saike.zwx,zhaoyiran.zyr,l.bing}@alibaba-inc.com anhtuan.luu@ntu.edu.sg
@@ 1 Introduction
Large language models (LLMs) frequently demonstrate the capability to understand and generate text across multiple languages, a skill attributed to their training on vast corpora composed of texts from various languages OpenAI (2023); Shi et al. (2022); Muennighoff et al. (2023); Jiang et al. (2023); Nguyen et al. (2023). However, these datasets are often disproportionately dominated by English content Brown et al. (2020); Chowdhery et al. (2022); Workshop et al. (2023), resulting in an English-centric bias in LLMs. This imbalance can subsequently hinder the models’ proficiency in other languages, often leading to suboptimal performance in non-English contexts Ahuja et al. (2023); Lai et al. (2023); Zhang et al. (2023b).
@@
To enhance performances in multilingual natural language processing (NLP) tasks with English-centric language models, translating training or test data into English has proven an effective strategy Conneau et al. (2018); Ponti et al. (2020); Artetxe et al. (2023); Moghe et al. (2023). Recent investigations have expanded this idea by incorporating translation, either implicitly or explicitly, into the intermediate stages of prompting LLMs Huang et al. (2023); Qin et al. (2023b); Etxaniz et al. (2023) for multilingual NLP tasks. For example, Shi et al., 2022 demonstrates that translating test questions into English enhances performance on multilingual reasoning tasks, as illustrated in Figure 2(a). Similarly, Huang et al., 2023 and Etxaniz et al., 2023 have shown that prompting LLMs to first translate or comprehend questions in English, then solve them step by step, improves performance.
@@
Despite these advancements, methodologies in various studies differ significantly, and the impact of translation on multilingual task performance remains underexplored. Furthermore, these studies focus on specific NLP tasks and English-centric LLMs, with less attention given to real-world user requests in various languages. This gap highlights a need for more nuanced research into the effectiveness of translation techniques across multilingual contexts. As shown in Figure 1, we hypothesize that English-centric LLMs generally perform better with English translations of prompts, while "Multilingual LLMs" excel with native prompts, particularly for tasks highly dependent on language.
In this study, we perform an in-depth analysis of the utility of translation with large language models for various scenarios. Firstly, we compare translating multilingual tasks into English, with an optional step of translating responses back into the original languages (i.e., the “translate-test” method), against several baselines on multilingual NLP tasks. Secondly, we extend the evaluation to real user queries, which are more likely to contain knowledge related to culture and language. Thirdly, We broaden the scope of LLM evaluations to include non-English-centric models to explore how they differ in behavior from English-centric LLMs.
@@
The main contributions of this work can be summarized as follows:
  • We compare various multilingual prompting strategies in NLP tasks, finding that translation remains a strong baseline even for LLMs. Our study also identifies factors that affect performance in multilingual tasks.
  • We broaden our multilingual evaluation to encompass actual user queries and LLMs that are not primarily centered on English, thereby addressing the limitations of previous works.
  • Our study reveals the shortcomings of current multilingual evaluations, underscoring the need for more comprehensive assessments that broaden the range of benchmarks and diversify the types of LLMs considered.
@@ 2.1.2 Models
This is the most capable and cost-effective model in the GPT-3.5222https://platform.openai.com/docs/models/gpt-3-5 family optimized for chat. We chose the latest version (gpt-3.5-turbo-1106) when starting the experiment.
@@ 2.1.3 Prompting Strategies
Considering the language of instruction, the use of chain-of-thought reasoning, and translation tools, we assess the following experimental strategies. We employ a zero-shot approach for all strategies, given that our selected models are fine-tuned to follow instructions.
@@
This is to directly pose the questions without employing additional prompting strategies such as chain-of-thought. Both the query information and the instruction provided are presented in the original language.
@@ 2.2 Main Results
The main results are shown in Table 1. We notice that Trans-Google, despite simple, demonstrates the highest overall performance across various models and tasks. While it may not always achieve top performance, it consistently delivers commendable results for both high and low-resource languages. Besides this, we can have the following observations: 1) Utilizing English instructions generally enhances performance across various tasks, regardless of the integration of chain-of-thought. This finding aligns with those reported by Lai et al., 2023. 2) chain-of-thought is quite helpful for strong LLMs like ChatGPT and reasoning tasks like MGSM. For weaker models and tasks that can be answered directly, the basic prompt may be a better option. 3) On average, EN-CoT underperforms compared to Trans-Google for both high and low-resource languages. While EN-CoT surpasses Trans-NLLB in high-resource languages, it falls short in low-resource ones. We hypothesize that this discrepancy arises because LLMs excel in high-resource languages but need external translation systems to handle low-resource languages effectively. More results and discussions are in Appendix A.1.4.
@@ 2.3 Analysis and Discussions
Table 2: Pearson correlation coefficient between MGSM accuracy and five language distances between English and that language. (*p < 0.05, two-tailed)
@@ 3 Translation for Real User Queries
The results of NLP tasks demonstrate that translating into English is a straightforward yet efficient method to address tasks in non-English languages. However, these benchmarks typically focus on specific linguistic aspects, which may not fully encapsulate the breadth and complexity of real-world user queries that often span an extensive range of topics and require nuanced comprehension. Moreover, these benchmarks are often constructed by translating the original English data into other languages Shi et al. (2022); Ponti et al. (2020); Conneau et al. (2018); Yang et al. (2019); Hasan et al. (2021). This approach leads to datasets that are not truly challenging, as they miss the rich culture-specific elements crucial for truly nuanced language understanding for different languages. To assess the impact of translation on real-world queries, we extract user requests from ShareGPT [^3] https://sharegpt.com/, a website to share real conversations with ChatGPT.
@@ 3.1 Experiment Setup
We selected 10 languages, ranging from high to low resource, and randomly sampled 100 requests for each language. However, for Romanian (ro), Ukrainian (uk), and Norwegian (no), we sampled 53, 98, and 53 requests respectively, due to the limited number of samples available from the source dataset. Since the queries can be in various formats, we only compare two promoting strategies: 1) original queries; and 2) translated queries with Google Translate API. For the second option, we translate the output back to the original language for consistency. To evaluate the quality of the responses, we use GPT-4o [^4] https://openai.com/index/hello-gpt-4o/(gpt-4o-2024-05-13) as the judge. The prompt for the judge is shown in Figure 8 in the Appendix, which is adapted from Zheng et al. (2023). With this prompt, each response will get a score from 1 to 10.
Figure 6: Results of four LLMs on M3Exam (a) language and (b) social science subject categories. Note that in M3Exam, not all subjects are available in every language, resulting in a discrepancy in the number of languages for the two subjects.
@@ 3.2 Main Results
To determine if answering user queries requires local cultural knowledge, we used GPT-4o with a specially crafted prompt to analyze queries in multiple languages (Figure 9 in the Appendix). Results in Table 14 in the Appendix show that 30% to 74% of queries per language require cultural knowledge, highlighting the rich cultural elements in the data. Further analysis of the ShareGPT subsets requiring local cultural knowledge is in Appendix A.2.
@@ 3.3 Analysis and Discussions
Based on the previous results, ChatGPT and Llama-2-70B-chat both tend to be English-centric but ChatGPT demonstrates certain behaviors of a "Multilingual LLM". An ideal "Multilingual LLM" should exhibit equivalent performance across all languages. Consequently, we broaden our analysis to include non-English-centric LLMs and assess their performance across various tasks.
@@
For the evaluation dataset, we choose M3Exam Zhang et al. (2023a), as the questions are real-world natural data from different languages instead of translating from English data and require strong multilingual proficiency and cultural knowledge to perform well. For example, the question about a Swahili proverb shown in Figure 2(b) requires local knowledge to answer correctly. We select the language and social science subject categories, which likely contain more native cultural knowledge, and evaluate up to 500 samples per language. For the prompting strategies, we follow (Zhang et al., 2023a) and use Monolingual, EN-Instruct and EN-Translation, which correspond to Native-Basic, EN-Basic  and Trans-Google in Section 2, for comparison.
Based on the results shown in Figure 6, we have the following observations: 1) For ChatGPT, translation may not always result in improved performance. This observation aligns with the conclusions drawn in the study by Zhang et al., 2023a. The effectiveness of translation largely depends on whether translation errors outweigh any potential gains in better comprehension. 2) Translation helps Llama-2-70B-chat in all the languages, suggesting that the model’s underperformance is due to poor language understanding rather than limitations of cultural knowledge. 3) Qwen1.5-72B-Chat and Yi-34B-Chat excel in Chinese proficiency. The translation process hurts Chinese performance, highlighting the significant influence of translationese on comprehension. Despite this, translation may still enhance performance in other languages, notably for Yi-34B-Chat, indicating that they are far from ideal multilingual LLMs.
@@
Prompt type MGSM XCOPA XNLI PAWS-X MKQA XL-Sum AVG
Qwen1.5-72B-Chat
Native-Basic 78.8 93.0 55.8 71.8 36.6 41.3 62.9
EN-Basic 77.2 97.0 73.0 73.0 32.7 39.7 65.4
Native-CoT 83.2 95.8 46.4 72.2 35.8 39.5 62.1
EN-CoT 81.6 97.2 71.2 70.6 34.9 38.6 65.7
XLT 78.4 97.8 77.4 67.6 20.8 35.3 62.9
Trans-Google 81.6 94.6 63.8 68.4 45.7 31.3 64.2
Trans-NLLB 58.8 88.2 61.4 70.4 32.0 28.5 56.5
Yi-34B-Chat
Native-Basic 63.2 92.6 46.0 43.6 13.4 36.9 49.3
EN-Basic 66.8 93.6 52.6 74.6 15.5 35.1 56.4
Native-CoT 65.2 91.8 42.6 43.6 13.0 36.6 48.8
EN-CoT 70.0 93.6 48.2 74.8 12.1 33.1 55.3
XLT 56.0 93.2 69.2 65.6 7.5 31.3 53.8
Trans-Google 71.2 94.0 49.6 70.8 24.5 36.3 57.7
Trans-NLLB 56.0 86.6 48.8 68.2 22.9 28.5 51.8
@@
To further understand the differences between English-centric LLMs and non-English-centric LLMs, we analyze the layerwise language distribution for Llama-2-7B-Chat and Qwen1.5-7B-Chat, using the method proposed by Zhao et al., 2024. We decode the embedding after each layer and identify each token into different languages with CLD3555https://github.com/google/cld3. As shown in Figure 7, the two LLMs process Chinese prompts differently. While the hidden representations of Qwen1.5-7B-Chat are mainly in Chinese, those of Llama-2-7B-Chat are in various other languages. We hypothesize that processing the information in native without conversion avoids the information loss, making it more suitable for processing multilingual tasks.
@@ 4 Related Work
The translate-test is a popular technique used to refine the performance of multilingual NLP benchmarks Conneau et al. (2018); Ponti et al. (2020); Artetxe et al. (2023); Moghe et al. (2023). In the era of LLMs, various strategies have been developed to enhance the performance of LLMs using multilingual datasets. Shi et al., 2022 discovered that EN-CoT outperforms Native-CoT. Huang et al., 2023 introduced cross-lingual-thought prompting to minimize language disparities. In parallel, Qin et al., 2023b introduced cross-lingual prompting, and Etxaniz et al., 2023 suggested self-translate to elevate their performances. Effective in translating prompts into English, these methods excel in NLP tasks but remain uncertain in real-world applications. Their success hinges on the English-centric nature of the LLMs. Our study evaluates translation effectiveness across NLP tasks, real user queries, and non-English-centric LLMs, revealing the limitations of these methods.

### v3

@@ Abstract
Large language models (LLMs) have demonstrated multilingual capabilities, yet they are mostly English-centric due to the imbalanced training corpora. While prior works have leveraged this bias to enhance multilingual performance through translation, they have been largely limited to natural language processing (NLP) tasks. In this work, we extend the evaluation to real-world user queries and non-English-centric LLMs, offering a broader examination of multilingual performance. Our key contribution lies in demonstrating that while translation into English can boost the performance of English-centric LLMs on NLP tasks, it is not universally optimal. For culture-related tasks that need deep language understanding, prompting in the native language proves more effective as it better captures the nuances of culture and language. Our experiments expose varied behaviors across LLMs and tasks in the multilingual context, underscoring the need for a more comprehensive approach to multilingual evaluation. Therefore, we call for greater efforts in developing and evaluating LLMs that go beyond English-centric paradigms. [^1] Our code is publicly available at https://github.com/DAMO-NLP-SG/translation-all-you-need.
@@ A Study on Solving Multilingual Tasks with Large Language Models
Chaoqun Liuthanks: Chaoqun Liu is under the Joint PhD Program between DAMO Academy and Nanyang Technological University.   12  Wenxuan Zhangthanks: Wenxuan Zhang is the corresponding author.  23  Yiran Zhao24  Anh Tuan Luu1  Lidong Bingthanks: Work done while at Alibaba Group.  5 1Nanyang Technological University, Singapore; 2DAMO Academy, Alibaba Group, Singapore; 3Hupan Lab, 310023, Hangzhou, China; 4National University of Singapore; 5Shanda AI Research Institute {chaoqun.liu,saike.zwx}@alibaba-inc.com; lidong.bing@shanda.com
@@ 1 Introduction
Large language models (LLMs) frequently demonstrate the capability to understand and generate text across multiple languages, a skill attributed to their training on vast corpora composed of texts from various languages OpenAI (2023); Shi et al. (2022); Muennighoff et al. (2023); Jiang et al. (2023); Nguyen et al. (2023). However, these datasets are often disproportionately dominated by English content Brown et al. (2020); Chowdhery et al. (2022); Workshop et al. (2023); Lin et al. (2022), resulting in an English-centric bias in LLMs. This imbalance can subsequently hinder the models’ proficiency in other languages, often leading to suboptimal performance in non-English contexts Ahuja et al. (2023); Lai et al. (2023); Zhang et al. (2023b).
@@
To enhance performances in multilingual natural language processing (NLP) tasks with English-centric language models, translating training or test data into English has proven an effective strategy Conneau et al. (2018); Ponti et al. (2020); Artetxe et al. (2023); Moghe et al. (2023); Bareiß et al. (2024). Recent investigations have expanded this idea by incorporating translation, either implicitly or explicitly, into the intermediate stages of prompting LLMs Huang et al. (2023); Qin et al. (2023b); Etxaniz et al. (2023) for multilingual NLP tasks. For example, Shi et al., 2022 demonstrates that translating test questions into English enhances performance on multilingual reasoning tasks, as illustrated in Figure 2(a). Similarly, Huang et al., 2023 and Etxaniz et al., 2023 have shown that prompting LLMs to first translate or comprehend questions in English, then solve them step by step, improves performance.
@@
Despite these advancements, methodologies in various studies differ significantly, and the impact of translation on multilingual task performance remains underexplored. Furthermore, these studies focus on specific NLP tasks and English-centric LLMs, but did not study real-world user queries in various languages. This gap highlights a need for more nuanced research into the effectiveness of translation techniques across multilingual contexts. As shown in Figure 1, we hypothesize that English-centric LLMs generally perform better with English translations of prompts, while "Multilingual LLMs" excel with native prompts, particularly for tasks highly dependent on language.
To address the limitations of existing empirical studies, we perform an in-depth analysis of the utility of translation with large language models for various scenarios. Firstly, we compare translating multilingual tasks into English, with an optional step of translating responses back into the original languages (i.e., the “translate-test” method), against several baselines on multilingual NLP tasks. Secondly, we extend the evaluation to real user queries, which are more likely to contain knowledge related to culture and language. Thirdly, we broaden the scope of LLM evaluations to include non-English-centric models to explore how they differ in behavior from English-centric LLMs. To the best of our knowledge, this is the first work to analyze the impacts of translating real user queries on multilingual LLMs.
@@
The main contributions of this work are:
  • We conduct a comprehensive comparison of multilingual prompting strategies in NLP tasks, finding that translation remains a strong baseline even for LLMs, and identifying factors impacting multilingual performance.
  • We expand multilingual evaluation to include actual user queries and and non-English-centric LLMs, addressing the limitations of previous studies.
  • We expose critical gaps in current multilingual evaluations, underscoring the need for more comprehensive benchmarks and a broader range of LLMs.
@@ 2.1.2 Models
This is the most capable and cost-effective model in the GPT-3.5333https://platform.openai.com/docs/models/gpt-3-5 family optimized for chat. We chose the latest version (gpt-3.5-turbo-1106) for the experiment.
@@ 2.1.3 Prompting Strategies
We assess experimental strategies based on language of instruction, chain-of-thought reasoning, and translation tools, using a zero-shot approach as the selected models are fine-tuned for instruction-following.
@@
The questions are posed directly without using prompting strategies like chain-of-thought. Both the query and instructions are presented in their original language.
@@ 2.2 Main Results
The main results are shown in Table 1. We notice that Trans-Google, despite simple, demonstrates the highest overall performance across various models and tasks. While it may not always achieve top performance, it consistently delivers commendable results for both high and low-resource languages. Besides this, we can have the following observations: 1) Utilizing English instructions generally enhances performance across various tasks, regardless of the integration of chain-of-thought. This finding aligns with those reported by Lai et al., 2023. 2) chain-of-thought is quite helpful for strong LLMs like ChatGPT and reasoning tasks like MGSM. For weaker models and tasks that can be answered directly, the basic prompt may be a better option. 3) On average, EN-CoT underperforms compared to Trans-Google for both high and low-resource languages. While EN-CoT surpasses Trans-NLLB in high-resource languages, it falls short in low-resource ones. We hypothesize that this discrepancy arises because LLMs excel in high-resource languages but need external translation systems to handle low-resource languages effectively.
These findings are also applicable to smaller models, such as Mistral-7B-Instruct, as demonstrated in Table 6 in the Appendix. This suggests that the observations generalize well across different model types and sizes. Further results and discussions are provided in Appendix A.1.4.
@@ 2.3 Analysis and Discussions
Table 2: Pearson correlation coefficient between MGSM accuracy and five language distances between English and that language. A lower value indicates higher correlation due to the negative coefficients.(*p < 0.05, two-tailed)
@@ 3 Translation for Real User Queries
NLP tasks typically focus on specific linguistic aspects, which may not fully encapsulate the breadth and complexity of real-world user queries which cover diverse topics and require nuanced comprehension. Moreover, these benchmarks are often constructed by translating from the English data Shi et al. (2022); Ponti et al. (2020); Conneau et al. (2018); Yang et al. (2019); Hasan et al. (2021). This approach leads to datasets that are not truly challenging, as they miss the rich culture-specific elements crucial for truly nuanced language understanding for different languages. To assess the impact of translation on real-world queries, we extract user requests from ShareGPT [^4] https://sharegpt.com/, a website to share real conversations with ChatGPT.
@@ 3.1 Experiment Setup
We selected 10 languages, ranging from high to low resource, and randomly sampled 100 requests for each language. However, for Romanian (ro), Ukrainian (uk), and Norwegian (no), we sampled 53, 98, and 53 requests respectively, due to the limited number of samples available from the source dataset. Since the queries can be in various formats, we only compare two prompting strategies: 1) original queries; and 2) translated queries with Google Translate API. For the second option, we translate the output back to the original language for consistency. To evaluate the quality of the responses, we use GPT-4o [^5] https://openai.com/index/hello-gpt-4o/(gpt-4o-2024-05-13) as the judge. The prompt for the judge is shown in Figure 8 in the Appendix, which is adapted from Zheng et al. (2023). With this prompt, each response will get a score from 1 to 10.
Figure 6: Accuracies of four LLMs on M3Exam (a) language and (b) social science subject categories. In M3Exam, not all subjects are available in every language, causing a difference in language coverage between the two subjects.
@@ 3.2 Main Results
To determine if answering user queries requires local cultural knowledge, we used GPT-4o with a specially crafted prompt to analyze queries in multiple languages (Figure 9 in the Appendix). Results in Table 14 in the Appendix show that 30% to 74% of queries per language require cultural knowledge, highlighting the rich cultural elements in the data. Further analysis of the ShareGPT subsets requiring local cultural knowledge is in Appendix A.2. We also conduct additional experiments, detailed in Appendix A.2.1, to verify that advanced LLMs can reliably assess the quality of responses.
@@ 3.3 Analysis and Discussions
Based on the previous results, ChatGPT and Llama-2-70B-chat both tend to be English-centric but ChatGPT demonstrates certain behaviors of a "Multilingual LLM". Consequently, we broaden our analysis to include non-English-centric LLMs and assess their performance across various tasks.
@@
For the evaluation dataset, we choose M3Exam Zhang et al. (2023a), as the questions are real-world natural data from different languages instead of translating from English and require strong multilingual proficiency and cultural knowledge to perform well. For example, the question about a Swahili proverb in Figure 2(b) requires local knowledge to answer correctly. We select the language and social science subject categories, which likely contain more native cultural knowledge, and evaluate up to 500 samples per language.
Based on the results shown in Figure 6, we have the following observations: 1) For ChatGPT, translation may not always result in improved performance. This observation aligns with the conclusions in the study by Zhang et al., 2023a. The effectiveness of translation largely depends on whether translation errors outweigh any potential gains in better comprehension. 2) Translation helps Llama-2-70B-chat in all the languages, suggesting that the model’s underperformance is due to poor language understanding rather than limitations of cultural knowledge. 3) Qwen1.5-72B-Chat and Yi-34B-Chat excel in Chinese proficiency. The translation hurts Chinese performance, highlighting the significant influence of translationese on comprehension. Despite this, it may boost performance in other languages, notably for Yi-34B-Chat, indicating that they are far from ideal multilingual LLMs.
@@
Prompt type Qwen1.5-72B-Chat Yi-34B-Chat
MGSM XCOPA XNLI PAWS-X MKQA XL-Sum AVG MGSM XCOPA XNLI PAWS-X MKQA XL-Sum AVG
Native-Basic 78.8 93.0 55.8 71.8 36.6 41.3 62.9 63.2 92.6 46.0 43.6 13.4 36.9 49.3
EN-Basic 77.2 97.0 73.0 73.0 32.7 39.7 65.4 66.8 93.6 52.6 74.6 15.5 35.1 56.4
Native-CoT 83.2 95.8 46.4 72.2 35.8 39.5 62.1 65.2 91.8 42.6 43.6 13.0 36.6 48.8
EN-CoT 81.6 97.2 71.2 70.6 34.9 38.6 65.7 70.0 93.6 48.2 74.8 12.1 33.1 55.3
XLT 78.4 97.8 77.4 67.6 20.8 35.3 62.9 56.0 93.2 69.2 65.6 7.5 31.3 53.8
Trans-Google 81.6 94.6 63.8 68.4 45.7 31.3 64.2 71.2 94.0 49.6 70.8 24.5 36.3 57.7
Trans-NLLB 58.8 88.2 61.4 70.4 32.0 28.5 56.5 56.0 86.6 48.8 68.2 22.9 28.5 51.8
@@
To further understand the differences between English-centric LLMs and non-English-centric LLMs, we analyze the layerwise language distribution for Llama-2-7B-Chat and Qwen1.5-7B-Chat, using the method proposed by Zhao et al., 2024. We decode the embedding after each layer and identify each token into different languages with CLD3666https://github.com/google/cld3. As shown in Figure 7, the two LLMs process Chinese prompts differently. While the hidden representations of Qwen1.5-7B-Chat are mainly in Chinese, those of Llama-2-7B-Chat are in various other languages. We hypothesize that processing the information in native without conversion avoids the information loss, making it more suitable for processing multilingual tasks. In addition, we examine the layerwise language distribution in larger models, specifically Llama-2-70B-Chat and Qwen1.5-72B-Chat, as shown in Figure 12 within Appendix A.3.
@@ 4 Related Work
The translate-test is a popular technique used to refine the performance of multilingual NLP benchmarks Conneau et al. (2018); Ponti et al. (2020); Artetxe et al. (2023); Moghe et al. (2023); Qi et al. (2022); Huang et al. (2022). In the era of LLMs, various strategies have been developed to enhance the performance of LLMs using multilingual datasets. Shi et al., 2022 discovered that EN-CoT outperforms Native-CoT. Huang et al., 2023 introduced cross-lingual-thought prompting to minimize language disparities. In parallel, Qin et al., 2023b introduced cross-lingual prompting, and Etxaniz et al., 2023 suggested self-translate to elevate their performances. Effective in translating prompts into English, these methods excel in NLP tasks but remain uncertain in real-world applications. Their success hinges on the English-centric nature of the LLMs. Our study evaluates translation effectiveness across NLP tasks, real user queries, and non-English-centric LLMs, revealing the limitations of these methods.